# N2: A Unified Python Package and Test Bench for Nearest Neighbor-Based Matrix Completion

### v1

@@ A.1 Factor model
For the scalar matrix completion problem, i.e., 1 with $n=1$, the main goal is to learn (or impute) the mean of the underlying distribution $\theta_{i,t}$ for any missing entries [37, 21, 20, 4, 48, 49]. The majority of this literature assumes (i) an additive noise model $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ for centered i.i.d. sub-Gaussian noise $\varepsilon$ and (ii) mean factor model, i.e., $\theta_{i,t}=f(u_{i},v_{t})$ for some latent factors $u_{i},v_{t}$ and real valued function $f$.
[…]
@@ A.2 Missingness pattern
For both the scalar and distributional matrix completion problem 1, the missing pattern (i.e., how the missingness $A_{j,s}$ was generated) can be categorized into three classes using the taxonomy of [36]: missing-completely-at-random (MCAR), missing-at-random (MAR) and missing-not-at-random (MNAR). MCAR assumes that the missingness $A_{i,t}$ is exogenous (independently generated from all the randomness in the model) and i.i.d. with propensity ${\mathbb{P}}(A_{i,t}=1)=p>0$ for all $(i,t)$. MAR is a more challenging scenario compared to MCAR as missingness is not exogenous, but its randomness depends on the observations. Further, propensities $p_{i,t}$ may differ for entries $(i,t)$ but positivity still holds, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}>0$. An important instance for MAR is the adaptive randomized policies [22]. The MNAR setup is the most challenging as it assumes the missingness depends on the unobserved latent confounders, while positivity may also be violated, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}=0$. The staggered adoption pattern, where a unit remains treated once a unit is treated at some adoption time, is a popular example of MNAR, mainly because positivity is violated. See [2, 5] for more details on staggered adoption.
We briefly outline the structural assumptions existing nearest neighbor methods were shown to work with provable guarantees; for all the existing methods, factor models (with slightly different details; compare the mean factorization [37] and the distribution factorization [16, 23]) are all commonly assumed.
  • (Scalar matrix completion) The vanilla versions of nearest neighbors ($\mathsf{RowNN}$) in [37, 20] are shown to work for MCAR and MAR setup; the latter shows that simple nearest neighbors can provably impute the mean when the missingness is fully adaptive across all users and history. The variants of vanilla nearest neighbors $\mathsf{DRNN}$ [21] is proven to work under MCAR, while $\mathsf{TSNN}$ [48] is proven to work under unobserved confounding, i.e., MNAR.
  • (Distributional matrix completion) The $\mathsf{KernelNN}$ [16] is shown to recover the underlying distribution under MNAR, whereas $\mathsf{W_{2}NN}$ [23] is shown to work under MCAR.
@@ Appendix B Nearest neighbor algorithms
The nearest neighbor methods introduced in Tab. 1 are elaborated in this section. We present two versions of each method; the first version explicitly constructs neighborhoods instead of subtly embedding them in the weights $\mathcal{W}$ of the Average module, and the second version specifies how each methods can be recovered by applying the two modules, Distance and Average, sequentially.
@@ B.1 Vanilla nearest neighbors
We elaborate on the discussion in Rem. 1 and provide here a detailed algorithm based on the explicit construction of neighborhoods, which is essentially equivalent to $\mathsf{RowNN}$ in Tab. 1. The inputs are measurements $\mathcal{Z}$, missingness $\mathcal{A}$, the target index $(i,t)$, and the radius $\eta$.
@@
In practice, the input $\eta$ for $\mathsf{RowNN}$ should be optimized via cross-validation; we refer the reader to App. C for a detailed implementation.
@@
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@ 2 Initialize hyper-parameter ${\eta}\leftarrow(\eta_{1},0)$
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$ return $\widehat{\theta}_{i,t}$
@@ B.2 Two-sided and doubly-robust nearest neighbors
We elaborate on the variants of the vanilla nearest neighbors algorithm $\mathsf{TSNN}$ and $\mathsf{DRNN}$ in Tab. 1; we first elaborate on an equivalent version of each of the methods which explicitly constructs neighborhoods.
@@
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@ 2 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},\eta_{2})$
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$ return $\widehat{\theta}_{i,t}$
@@
For $\mathsf{DRNN}$ algorithm below, we consider $\mathcal{Z}$ and $\mathcal{A}$ to be $N\times T$ sized matrices, so that their transpose is well defined. Then note that $\mathsf{ColNN}$ is simply applying Alg. 1 with transposed observation matrices.
@@
1 Initialize $\mathsf{RowNN}\leftarrow$ Alg. 1 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
2 Initialize $\mathsf{ColNN}\leftarrow$ Alg. 1 with input $(\mathcal{Z}^{T},\mathcal{A}^{T},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
3 Initialize $\mathsf{TSNN}\leftarrow$ Alg. 2 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and ${\eta}\leftarrow(\eta_{1},\eta_{2})$
Calculate $\widehat{\theta}_{i,t}\leftarrow\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$ return $\widehat{\theta}_{i,t}$
@@ B.3 Distributional nearest neighbors
Unlike the scalar nearest neighbor methods, distributional nearest neighbors necessitate a distributional notion of distance between rows and columns of matrix and a distributional analog of averaging. [16] and [23] use maximum mean discrepency (in short $\mathsf{MMD}$) of kernel mean embeddings [40] and Wasserstein metric (in short $\mathsf{W}_{2}$[10] respectively both for defining the distance between rows / columns and for averaging. The corresponding barycenters of $\mathsf{MMD}$ and $\mathsf{W}_{2}$ [14, 9] are used for averaging, and so the methods are coined kernel nearest neighbors (in short $\mathsf{KernelNN}$) and Wasserstein nearest neighbors (in short $\mathsf{W_{2}NN}$) respectively.
@@
    $\widehat{\mu}^{\mathsf{W}_{2}}_{i,t,\eta}$ $:=\mathop{\mathrm{missing}}{argmin}_{\mu}\sum_{j\in\in\mathbf{N}_{t,\eta}^{\mathsf{W}_{2}}}\mathsf{W}_{2}^{2}(\mu,\mu_{j,t}^{Z}).$ (16)
For further details on the $\mathsf{W}_{2}$ and $\mathsf{MMD}$ algorithms see [23] and [16], respectively.
@@
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow$ $\widehat{\mathsf{MMD}}^{2}_{\mathbf{k}}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$ or $\widehat{\mathsf{W}}_{2}^{2}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$
@@ 3 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},0)$
Calculate average $\widehat{\mu}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$ return $\widehat{\mu}_{i,t}$
@@ B.4 Adaptively weighted nearest neighbors
We elaborate on the adaptive variant of the vanilla nearest neighbor algorithm $\mathsf{AWNN}$ as mentioned in Sec. 2.2 and Tab. 1. The input are measurements $\mathcal{Z}$, and missingness $\mathcal{A}$. Note that there is no need for radius parameter $\eta$ and hence no CV.
@@
The optimization problem in 19 can be solved exactly in linear time (worst case complexity) using convex optimization [49]. $\mathsf{AWNN}$ doesn’t rely on radius parameter $\eta$. Not only it automatically assigns neighbors to ${(i,t)^{th}}$ entry during its weight calculation(non-neighbors get zero weight), but also takes into account the distance of the neighbors from the ${(i,t)^{th}}$ entry. The closer neighors get higher weights and vice - versa.
@@
1 Initialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@
Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$ return $\widehat{\theta}_{i,t}$
@@ Appendix C Cross-Validation
For each nearest neighbor method, we use cross-validation to optimize hyperparameters including distance thresholds and weights, depending on which nearest neighbor algorithm is chosen. Specifically, for each experiment, we choose a subset of the training test to optimize hyperparameters by masking those matrix cells and then estimating the masked values. We utilize the HyperOpt library [13] to optimize (possibly multiple) hyperparamters using the Tree of Parzen Estimator [7], a Bayesian optimization method. Our package supports both regular distance thresholds and percentile-based thresholds, which adapt to the distances calculated within the specific dataset.
@@ Appendix D Case Study Details
The boxplots are generated using matplotlib’s [30] standard boxplot function. The box shows the first, second, and third quartiles. The bottom line shows the first quartile minus the 1.5$\times$ the interquartile range. The top line shows the third quartile plus 1.5$\times$ the interquartile range. All experiments are run on standard computing hardware (MacBook Pro with an M2 Pro CPU with 32 GB of RAM).
@@ D.1 Synthetic data generation
Generate $Z_{i,t}=X_{i,t}\sim N(\theta_{i,t},\sigma^{2})$, i.e., scalar matrix completion setting, with a linear factor structure $\theta_{i,t}=u_{i}v_{t}$. Row latent factors $u_{i}\in^{4}$ are i.i.d. generated across $i=1,...,N$, where each entry of $u_{i}$ follow a uniform distribution with support $[-0.5,0.5]$; column latent factors $v_{t}\in^{4}$ are generated in an identical manner. The missingness is MCAR with propensity $p_{i,t}=0.5$ for all $i$ and $t$. Further, the size of column and rows are identical $N=T$. For the left panel in Fig. 1, the noise level is set as $\sigma=0.001$ and for the right panel $\sigma=1$.
@@ D.2 HeartSteps V1
The mobile application was designed to send notifications to users at various times during the day to encourage anti-sedentary activity such as stretching or walking. Participants could be marked as unavailable during decision points if they were in transit or snoozed their notifications, so notifications were only sent randomly if a participant was available and were never sent if they were unavailable. To process the data in the framework of 1, we let matrix entry $Z_{i,t}$ be the average one hour step count for participant $i$ and decision point $t$ when a notification is sent (i.e. $A_{i,t}=1$) and unknown when a notification is not sent (i.e. $A_{i,t}=0$). The treatment assignment pattern is represented as the 37 x 200 matrix visualized in Fig. 6. We use the dataset downloaded from https://github.com/klasnja/HeartStepsV1 (CC-BY-4.0 License).

### v2

@@ A.1 Factor model
For the scalar matrix completion problem, i.e., ˜2 with $n=1$, the main goal is to learn (or impute) the mean of the underlying distribution $\theta_{i,t}$ for any missing entries [li2019nearest, dwivedi2022doubly, dwivedi2022counterfactual, agarwal2023causal, sadhukhan2024adaptivity, sadhukhan2025adaptivelyweightednearestneighborsmatrix]. The majority of this literature assumes (i) an additive noise model $\mu_{i,t}=\theta_{i,t}+\varepsilon_{i,t}$ for centered i.i.d. sub-Gaussian noise $\varepsilon$ and (ii) mean factor model, i.e., $\theta_{i,t}=f(u_{i},v_{t})$ for some latent factors $u_{i},v_{t}$ and real valued function $f$.
For the distributional matrix completion problem (i.e., ˜2 with $n>1$) the main goal is to learn the underlying distribution itself [choi2024learning, feitelberg2024distributional]; a factor model is imposed on the distribution as a whole. For instance, a factor model is assumed on the kernel mean embedding of distributions; that is, there exist latent factors $u_{i}$ and $v_{t}$ and an operator $g$ such that $\int\mathbf{k}(x,\cdot)d\mu_{i,t}(x)=g(u_{i},v_{t})$.
@@ A.2 Missingness pattern
For both the scalar and distributional matrix completion problem ˜2, the missing pattern (i.e., how the missingness $A_{j,s}$ was generated) can be categorized into three classes using the taxonomy of [little2019statistical]: missing-completely-at-random (MCAR), missing-at-random (MAR) and missing-not-at-random (MNAR). MCAR assumes that the missingness $A_{i,t}$ is exogenous (independently generated from all the randomness in the model) and i.i.d. with propensity ${\mathbb{P}}(A_{i,t}=1)=p>0$ for all $(i,t)$. MAR is a more challenging scenario compared to MCAR as missingness is not exogenous, but its randomness depends on the observations. Further, propensities $p_{i,t}$ may differ for entries $(i,t)$ but positivity still holds, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}>0$. An important instance for MAR is the adaptive randomized policies [dwivedi2022personalization]. The MNAR setup is the most challenging as it assumes the missingness depends on the unobserved latent confounders, while positivity may also be violated, i.e., $\min_{i\in[N],t\in[T]}p_{i,t}=0$. The staggered adoption pattern, where a unit remains treated once a unit is treated at some adoption time, is a popular example of MNAR, mainly because positivity is violated. See [athey2021matrix, athey2022design] for more details on staggered adoption.
We briefly outline the structural assumptions existing nearest neighbor methods were shown to work with provable guarantees; for all the existing methods, factor models (with slightly different details; compare the mean factorization [li2019nearest] and the distribution factorization [choi2024learning, feitelberg2024distributional]) are all commonly assumed.
  • (Scalar matrix completion) The vanilla versions of nearest neighbors ($\mathsf{RowNN}$) in [li2019nearest, dwivedi2022counterfactual] are shown to work for MCAR and MAR setup; the latter shows that simple nearest neighbors can provably impute the mean when the missingness is fully adaptive across all users and history. The variants of vanilla nearest neighbors $\mathsf{DRNN}$ [dwivedi2022doubly] is proven to work under MCAR, while $\mathsf{TSNN}$ [sadhukhan2024adaptivity] is proven to work under unobserved confounding, i.e., MNAR.
  • (Distributional matrix completion) The $\mathsf{KernelNN}$ [choi2024learning] is shown to recover the underlying distribution under MNAR, whereas $\mathsf{W_{2}NN}$ [feitelberg2024distributional] is shown to work under MCAR.
@@ Appendix B Nearest neighbor algorithms
The nearest neighbor methods introduced in Tab.˜1 are elaborated in this section. We present two versions of each method; the first version explicitly constructs neighborhoods instead of subtly embedding them in the weights $\mathcal{W}$ of the Average module, and the second version specifies how each methods can be recovered by applying the two modules, Distance and Average, sequentially.
@@ B.1 Vanilla nearest neighbors
We elaborate on the discussion in Rem.˜1 and provide here a detailed algorithm based on the explicit construction of neighborhoods, which is essentially equivalent to $\mathsf{RowNN}$ in Tab.˜1. The inputs are measurements $\mathcal{Z}$, missingness $\mathcal{A}$, the target index $(i,t)$, and the radius $\eta$.
@@
In practice, the input $\eta$ for $\mathsf{RowNN}$ should be optimized via cross-validation; we refer the reader to App.˜C for a detailed implementation.
@@
1
21exInitialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@ 3 Initialize hyper-parameter ${\eta}\leftarrow(\eta_{1},0)$
6 Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$
1exreturn $\widehat{\theta}_{i,t}$
@@ B.2 Two-sided and doubly-robust nearest neighbors
We elaborate on the variants of the vanilla nearest neighbors algorithm $\mathsf{TSNN}$ and $\mathsf{DRNN}$ in Tab.˜1; we first elaborate on an equivalent version of each of the methods which explicitly constructs neighborhoods.
@@
1
21exInitialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@ 3 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},\eta_{2})$
6 Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$
1exreturn $\widehat{\theta}_{i,t}$
@@
For $\mathsf{DRNN}$ algorithm below, we consider $\mathcal{Z}$ and $\mathcal{A}$ to be $N\times T$ sized matrices, so that their transpose is well defined. Then note that $\mathsf{ColNN}$ is simply applying Alg.˜1 with transposed observation matrices.
@@
1
21exInitialize $\mathsf{RowNN}\leftarrow$ Alg.˜1 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
3 Initialize $\mathsf{ColNN}\leftarrow$ Alg.˜1 with input $(\mathcal{Z}^{T},\mathcal{A}^{T},\eta,(i,t))$ and $\eta\leftarrow(\eta_{1},0)$
4 Initialize $\mathsf{TSNN}\leftarrow$ Alg.˜2 with inputs $(\mathcal{Z},\mathcal{A},\eta,(i,t))$ and ${\eta}\leftarrow(\eta_{1},\eta_{2})$
5 Calculate $\widehat{\theta}_{i,t}\leftarrow\mathsf{RowNN}+\mathsf{ColNN}-\mathsf{TSNN}$
1exreturn $\widehat{\theta}_{i,t}$
@@ B.3 Distributional nearest neighbors
Unlike the scalar nearest neighbor methods, distributional nearest neighbors necessitate a distributional notion of distance between rows and columns of matrix and a distributional analog of averaging. [choi2024learning] and [feitelberg2024distributional] use maximum mean discrepency (in short $\mathsf{MMD}$) of kernel mean embeddings [muandet2017kernel] and Wasserstein metric (in short $\mathsf{W}_{2}$[bigot2020] respectively both for defining the distance between rows / columns and for averaging. The corresponding barycenters of $\mathsf{MMD}$ and $\mathsf{W}_{2}$ [cohen2020estimating, bigot2017geodesic] are used for averaging, and so the methods are coined kernel nearest neighbors (in short $\mathsf{KernelNN}$) and Wasserstein nearest neighbors (in short $\mathsf{W_{2}NN}$) respectively.
@@
    $\widehat{\mu}^{\mathsf{W}_{2}}_{i,t,\eta}$ $:=\mathop{\mathrm{argmin}}_{\mu}\sum_{j\in\in\mathbf{N}_{t,\eta}^{\mathsf{W}_{2}}}\mathsf{W}_{2}^{2}(\mu,\mu_{j,t}^{Z}).$ (17)
For further details on the $\mathsf{W}_{2}$ and $\mathsf{MMD}$ algorithms see [feitelberg2024distributional] and [choi2024learning], respectively.
@@
1
21exInitialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow$ $\widehat{\mathsf{MMD}}^{2}_{\mathbf{k}}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$ or $\widehat{\mathsf{W}}_{2}^{2}(Z_{j,s},Z_{j^{\prime},s^{\prime}})$
@@ 4 Initialize tuning parameter ${\eta}\leftarrow(\eta_{1},0)$
7 Calculate average $\widehat{\mu}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},\mathcal{Z},\mathcal{A})$
1exreturn $\widehat{\mu}_{i,t}$
@@ B.4 Adaptively weighted nearest neighbors
We elaborate on the adaptive variant of the vanilla nearest neighbor algorithm $\mathsf{AWNN}$ as mentioned in Sec.˜2.2 and Tab.˜1. The input are measurements $\mathcal{Z}$, and missingness $\mathcal{A}$. Note that there is no need for radius parameter $\eta$ and hence no CV.
@@
The optimization problem in ˜20 can be solved exactly in linear time (worst case complexity) using convex optimization [sadhukhan2025adaptivelyweightednearestneighborsmatrix]. $\mathsf{AWNN}$ doesn’t rely on radius parameter $\eta$. Not only it automatically assigns neighbors to $(i,t)^{th}$} entry during its weight calculation(non-neighbors get zero weight), but also takes into account the distance of the neighbors from the $(i,t)^{th}$} entry. The closer neighors get higher weights and vice - versa.
@@
1
21exInitialize entry-wise metric $\widehat{\varphi}(Z_{j,s},Z_{j^{\prime},s^{\prime}})\leftarrow(Z_{j,s}-Z_{j^{\prime},s^{\prime}})^{2}$ and metric $\varphi(x,y)\leftarrow(x-y)^{2}$
@@
6 Calculate average $\widehat{\theta}_{i,t}\leftarrow\textsc{Average}(\varphi,\mathcal{W},Z,A)$
1exreturn $\widehat{\theta}_{i,t}$
@@ Appendix C Cross-Validation
For each nearest neighbor method, we use cross-validation to optimize hyperparameters including distance thresholds and weights, depending on which nearest neighbor algorithm is chosen. Specifically, for each experiment, we choose a subset of the training test to optimize hyperparameters by masking those matrix cells and then estimating the masked values. We utilize the HyperOpt library [bergstra2013making] to optimize (possibly multiple) hyperparamters using the Tree of Parzen Estimator [bergstra2011algorithms], a Bayesian optimization method. Our package supports both regular distance thresholds and percentile-based thresholds, which adapt to the distances calculated within the specific dataset.
@@ Appendix D Case Study Details
[…]
@@ D.1 Synthetic data generation
Generate $Z_{i,t}=X_{i,t}\sim N(\theta_{i,t},\sigma^{2})$, i.e., scalar matrix completion setting, with a linear factor structure $\theta_{i,t}=u_{i}v_{t}$. Row latent factors $u_{i}\in^{4}$ are i.i.d. generated across i=1,,Ni=1,...,N, where each entry of $u_{i}$ follow a uniform distribution with support [0.5,0.5][-0.5,0.5]; column latent factors $v_{t}\in^{4}$ are generated in an identical manner. The missingness is MCAR with propensity $p_{i,t}=0.5$ for all $i$ and $t$. Further, the size of column and rows are identical $N=T$. For the left panel in Fig.˜1, the noise level is set as $\sigma=0.001$ and for the right panel $\sigma=1$.
@@ D.2 HeartSteps V1
The mobile application was designed to send notifications to users at various times during the day to encourage anti-sedentary activity such as stretching or walking. Participants could be marked as unavailable during decision points if they were in transit or snoozed their notifications, so notifications were only sent randomly if a participant was available and were never sent if they were unavailable. To process the data in the framework of ˜2, we let matrix entry $Z_{i,t}$ be the average one hour step count for participant $i$ and decision point $t$ when a notification is sent (i.e. $A_{i,t}=1$) and unknown when a notification is not sent (i.e. $A_{i,t}=0$). The treatment assignment pattern is represented as the 37 x 200 matrix visualized in Fig.˜6. We use the dataset downloaded from https://github.com/klasnja/HeartStepsV1 (CC-BY-4.0 License).